If the energy of an electron in hydrogen atom is given by expression -1312/n^2 kJ mol^-1, find the energy required to excite the electron from ground state to second orbit.

Grade:11

1 Answers

Arun
4 years ago
 
Energy of 1st orbit=13.12*10^5(let's say E)
Thus, energy of second orbit=3.28*1-^5(Because E/n^2,let's say E*)
Thus, enery needed to ake from 1st to 2nd orbit=Difference of the two=9.84*10^5 J/mol